Anyone willing to help me?
Solve for y. Shoe algebraic steps and check your answer.
1-2y = 6y-11

Answer by lynnlo(4176)

1-2Y=6Y-11
1+-2Y=-11+6Y
1+-2Y+-6=-11+6Y+-6Y
1+-8Y=-11+6Y+-6Y
1+-8Y=-11+0
1+-8Y=-11
1+-1+-8Y=11+-1
0+-8Y=-11+-1
-8Y=-11+-1
Y=1.5